OpenStudy (anonymous):

Factor. 2c^2+9c+9 Factor. 2c^2+9c+9 @Mathematics

OpenStudy (anonymous):

(2c+3)(c+3)

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

yw 2 * 9 = 18 6 * 3 = 18 and 6+3 = 9